According to the nutrition information given, one serving of total cereal contains 23 grams of carbohydrates. If one serving is 3/4 cup, what amount of carbohydrates is contained in 2 1/4 cups of total cereal ?
Answer by CubeyThePenguin(3113)

2 1/4 cups = 9/4 cups = 3 servings = 69 grams of carbohydrate